# InsBank: Evolving Instruction Subset for Ongoing Alignment

### v1

@@ Abstract
Large language models (LLMs) typically undergo instruction tuning to enhance alignment. Recent studies emphasize that quality and diversity of instruction data are more crucial than quantity, highlighting the need to select diverse, high-quality subsets to reduce training costs. However, how to evolve these selected subsets alongside the development of new instruction data remains insufficiently explored. To achieve LLMs’ ongoing alignment, we introduce Instruction Bank (InsBank), a continuously updated repository that integrates the latest valuable instruction data. We further propose Progressive Instruction Bank Evolution (PIBE), a novel framework designed to evolve InsBank effectively and efficiently over time. PIBE employs a gradual data selection strategy to maintain long-term efficiency, leveraging a representation-based diversity score to capture relationships between data points and retain historical information for comprehensive diversity evaluation. This also allows for flexible combination of diversity and quality scores during data selection and ranking. Extensive experiments demonstrate that PIBE significantly outperforms baselines in InsBank evolution and is able to extract budget-specific subsets, demonstrating its effectiveness and adaptability [^1] Our code has been released on https://github.com/jiayinlp/InsBank.
@@
Jiayi Shi1222Equal contributions.,      Yiwei Li1222Equal contributions.,      Shaoxiong Feng2,      Peiwen Yuan1,      Xinglin Wang1, Yueqi Zhang1,      Chuyi Tan1,      Boyuan Pan2,      Huan Ren2,      Yao Hu 2,      Kan Li1333Corresponding author., 1 School of Computer Science, Beijing Institute of Technology 2 Xiaohongshu Inc
@@ 1 Introduction
Specifically, with the continuous emergence of instruction datasets (The timeline of part instruction datasets is shown in Appendix A), it becomes necessary to regularly update the instruction subset to incorporate the latest advanced instruction data in order to ensure ongoing improvements in the alignment capabilities of LLMs. Simultaneously, the subset size must be controlled to avoid excessive growth that could lead to increased training costs. To address these practical challenges, we propose a novel concept termed InsBank (Instruction Bank). As shown in Figure 1, it is initially built by selecting current available instruction data. When new datasets are proposed, the bank evolves by selecting new data while phasing out an equivalent amount of older data, thereby maintaining an optimized instruction subset. The data in InsBank is also ranked, enabling users to extract smaller subsets tailored to specific training budgets efficiently.
The orderliness of InsBank is achieved through an overall score that combines individual quality and diversity scores. Quality scores can be obtained through manual or model annotation, but measuring diversity requires a global comparison between data, leading to significant storage and computational costs. During InsBank evolution, the impact of new data on the overall distribution necessitates continuous adjustment of each sample’s diversity score. A straightforward approach would be to re-select data from all available instruction data during each evolution iteration. However, the massive volume of instruction data (Qin et al., 2024a) and its rapid growth (Longpre et al., 2023; Wang et al., 2023; Xu et al., 2023) make this approach prohibitively expensive. Moreover, existing methods struggle to effectively represent and combine diversity and quality scores for ranking purposes.
@@
To address these challenges, we propose Progressive Instruction Bank Evolution (PIBE) for continuous and efficient selection of the optimal instruction subset. PIBE employs a gradual manner of selection to evolve InsBank, ensuring long-term efficiency. Unlike the naive approach, PIBE significantly reduces evolution costs by excluding previously filtered-out data and focusing only on newly proposed and current InsBank data. However, the absence of past data changes the distribution of candidates, making it critical to preserve historical distribution information during evolution. Existing diversity-driven data selection methods (Liu et al., 2024; Wu et al., 2023) fall into two main categories: k-nearest neighbor (k-NN) (Dong et al., 2011) and geometry-based coreset sampling (Guo et al., 2022). However, both of them rely solely on local information from a few neighboring points, making it difficult to record and utilize the rich information of previously eliminated data. Consequently, they cannot capture global relationships between points or provide robust individual diversity scores for effective ranking. Inspired by Affinity Propagation (Frey and Dueck, 2007), we frame InsBank data selection as an exemplar election process, where the representativeness of each data point is quantified through an iterative voting mechanism. The representativeness further serves as the individual diversity score, and the voting results are passed to the next iteration as historical information to preserve the distribution of absent data. Moreover, existing data selection methods either prioritize quality or diversity (Chen et al., 2024), or address them sequentially (Liu et al., 2024), failing to consider both aspects equally. In contrast, our diversity score integrates seamlessly with the quality score, enabling comprehensive and flexible instruction selection and InsBank ranking.
@@
We simulate the process of instruction set development with five datasets and perform InsBank evolution on them with PIBE. We evaluate the general instruction following capability of fine-tuned models on AlpacaEval (Li et al., 2023b), MT-Bench (Zheng et al., 2023) and IFEval (Zhou et al., 2023b). Experimental results show that PIBE outperforms the baselines and successfully evolves the instruction bank in parallel with the development of instruction sets. Besides, analysis on orderliness of InsBank indicates that users can flexibly select a smaller subset based on their budget. Ours contributions can be summarized as follows:
@@ 3 Progressive Instruction Bank Evolution
Figure 2: The detailed pipeline of PIBE. It consists of four core elements: the gradual manner of evolution, the flow of historical information across evolution rounds, individual representation scoring for diversity evaluation, and the integration of quality and diversity scores for data selection and ranking
@@ 3.1 Gradual Evolution Formulation
In this work, we propose the instruction subset evolution task to build the InsBank. Denoting current available instruction data as $\mathcal{X}_{0}$, the instruction bank $\mathcal{B}_{\pi}^{0,m}$ of size $m$ is initialized through data selection which can be presented as $\mathcal{B}_{\pi}^{0,m}=\pi(\mathcal{X}_{0})$. Then, when new instruction dataset $\mathcal{X}_{1}$ is proposed, $\mathcal{B}_{\pi}^{0,m}$ should evolve itself to adapt to changes in data distribution. The naive manner of InsBank evolution can be represented as $\mathcal{B}_{\pi}^{1,m}=\pi(\mathcal{X}_{0},\mathcal{X}_{1})$ which can be extended to $\mathcal{B}_{\pi}^{t+1,m}=\pi(\mathcal{X}_{0},...,\mathcal{X}_{t},\mathcal{X}_{t+1})$ for future evolution. However, this manner requires substantial storage and computational resources to calculate diversity scores as $t$ continues to increase. To improve the long-term evolution efficiency, we propose a gradual manner where only the newly proposed instruction data $\mathcal{X}_{t+1}$ along with the data participated in last round of evolution $\mathcal{X}_{t}+\mathcal{B}_{\pi}^{t-1,m}$ are involved into the current round of evolution, and the evolution can be represented as $\mathcal{B}_{\pi}^{t+1,m}=\pi(\mathcal{X}_{t+1},\mathcal{X}_{t}+\mathcal{B}_{t-1}^{m})$.
@@ 3.2 Historical Information Flowing
The structure of $M_{t}$ is depicted in Appendix E. The top-left part of $M_{t}$ contains responsibility values between data in $\mathcal{B}^{t,m}_{\pi}$, taken directly from $H_{t}$. The top-right part represents the suitability of newly proposed candidate data as exemplars for previously selected data, estimated using Eq. 4. Similarly, the bottom-left part represents the suitability of previously selected data as exemplars for newly proposed candidate data, estimated using Eq. 5. The bottom-right section is filled with the median values of the other three sections.
@@ 3.4 Integration of Diversity and Quality
In practice, we observe that further improving quality beyond a certain level can reduce the fine-tuned model’s performance. Additionally, when combining quality and diversity using linear methods, diversity scores often dominate the selection process. This occurs because quality, as a linear score, increases at a constant rate, even when excessively large values provide diminishing benefits. More details can be found in our experimental analysis of score combination (Section 4.4).
@@
Table 1: Comparison between different methods. For AlpacaEval and MT-Bench, we employ gpt-4o as annotator. The bold text indicates the best results, and the underlined text represents the second-best results.
To address this, we design a nonlinear mapping function for quality scores, shown in Eq. 11. Here, $Q_{p}$ denotes the $p$-th percentile, $r_{l}$ and $r_{h}$ represent the lower and upper percentiles, $S^{\prime}_{q}$ refers to the scaled quality scores, and $\sigma(\cdot)$ is the sigmoid function. The function, illustrated in Figure 7, leverages the sigmoid’s steepness in $(-2,2)$ to enhance the distinguishability of scores within $[\tau_{l},\tau_{h}]$, while flattening growth for scores above $\tau_{h}$. Data below $\tau_{l}$ are less considered, as such low-quality data are rarely selected into InsBank. Finally, we combine diversity with the nonlinear-mapped quality scores.
@@ 4.1 Experimental Setup
Training and Evaluation In this work, we fine-tune the Llama3-8B model (AI@Meta, 2024) on the selected InsBank unless otherwise specified. Following DEITA (Liu et al., 2024), we set the size of InsBank to 6k for the convenience of subset evolution. During training, we further restrict the trainable tokens and the number of conversation turns. We adopt AlpacaEval (Li et al., 2023b), MT-Bench (Zheng et al., 2023) and IFEval (Zhou et al., 2023b) for automatic model alignment performance evaluation. More details about training and evaluation can be found in Appendix B.
@@
  • DEITA Traverse the instruction pool in descending order of quality scores and involve the current sample to the selected subset if the largest cosine similarity between the current sample and the samples in the selected subset is less than the threshold (i.e. 0.9 following the raw setting of DEITA (Liu et al., 2024)).
@@ 4.4 Analysis
Effectiveness of Diversity and Quality
To validate the role of diversity in instruction data selection, we first construct a quality-controlled subset where all data have quality scores within the range of 4.5 to 5.0 (details in Appendix F). Using PIBE, we compute individual diversity scores for the subset, sort the data in descending order, and select the top 6k samples as the most diverse subset and the bottom 6k as the least diverse subset. The distributions of the two subsets are shown in Fig. 6. Before fine-tuning, we restrict the total trainable tokens to 2M. Results in Table 3 indicate that, with comparable quality, models trained on more diverse data achieve better performance.
@@
Table 3: The results of quality-controlled diversity effectiveness experiment. Here, Qua refers to the average quality score, and Div refers to the average diversity score.
When it comes to quality, the improvement from extremely low to high quality is clearly beneficial, as extremely low-quality subsets often contain noisy data, such as irrelevant or incomplete responses. However, is continuously improving quality always effective in the data selection process? To address this, we compare model performance fine-tuned on data selected by the following strategies in the final evolution iteration: (1) Diversity Greedy: selecting data with the highest diversity scores; (2) Quality Greedy: selecting data with the highest quality scores; and (3) PIBE. The results shown in Table 4 reveal a clear trade-off between diversity and quality. A purely greedy approach focusing on either aspect leads to suboptimal outcomes, while a balanced consideration of both proves more effective. This finding aligns with the main experiment results and suggests the existence of a balance point between diversity and quality, which we further investigate through combination methods in Section 4.4.
@@
Analysis of Score Combination
We experiment with the different combination methods to explore the contribution of quality and diversity in PIBE.
@@
Table 5: The results of different combination methods. Here, SP- refers to Spearman Correlation Coefficient, Diff refers to the difference value between SP-Qua and SP-Div.
We first explore the basic multiplication manner and the addition manner, and the results are reported in Table 5. Overall, regardless of whether addition or multiplication is used as the combination method, the experimental results exhibit a distinct trend of initially increasing and then decreasing as the influence of quality grows (i.e., with the increase of the $\gamma$ value). This finding supports the hypothesis that a balance point exists between diversity and quality.
@@ 5 Related Work
Quality-based Selection Humpback (Li et al., 2023a) selects high-quality samples through an iterative self-curation process where quality predictions are produced by the fine-tuned model of each turn. Recent works typically employ a GPT-model to annotate the data quality. For example, ALPAGASUS Chen et al. (2024) employs ChatGPT to score the accuracy of instruction data and select data according to a threshold.
@@
Model-specific Importance-based Selection Importance refers to the necessity of adding one sample into training set (Liu et al., 2024) whose indicator are typically model-specific (Xia et al., 2024; Li et al., 2024a). However, this work focuses on the general data selection and emphasizes the quality and diversity of selected data.
@@ 6 Conclusion
In this paper, we propose InsBank to address the ongoing challenge of evolving instruction datasets. PIBE integrates high-quality and representative data into InsBank, striking a balance between enhancing data diversity and quality, while maintaining long-term scalability and efficiency. By leveraging a representation-based diversity score with historical information, PIBE flexibly combines diversity and quality for data selection and ranking. Experimental results show PIBE outperforms baselines, providing more optimal and adaptable instruction subsets. The orderliness of InsBank also allows users to extract tailored subsets within budget constraints, supporting cost-effective training and the ongoing refinement of LLMs. This work paves the way for more dynamic and adaptable instruction tuning strategies, enhancing both the efficiency and effectiveness of LLM development over time.
@@
In this work, we focus on evaluating the diversity of individual instruction data and exploring the combination of diversity and quality scores. However, achieving a more precise assessment of data quality remains a valuable direction for future research. Additionally, due to the high memory cost of the Affinity Propagation algorithm, PIBE must process candidate data in batches to prevent memory overflow. Nevertheless, since the data selection time for each batch is short (typically around 60 seconds, depending on the number of convergence steps), such batch processing does not lead to a significant increase in PIBE’s time cost.

### v2

@@ Abstract
Large language models (LLMs) typically undergo instruction tuning to enhance alignment. Recent studies emphasize that quality and diversity of instruction data are more crucial than quantity, highlighting the need to select diverse, high-quality subsets to reduce training costs. However, how to evolve these selected subsets alongside the development of new instruction data remains insufficiently explored. To achieve LLMs’ ongoing alignment, we introduce Instruction Bank (InsBank), a continuously updated repository that integrates the latest valuable instruction data. We further propose Progressive Instruction Bank Evolution (PIBE), a novel framework designed to evolve InsBank effectively and efficiently over time. PIBE employs a gradual data selection strategy to maintain long-term efficiency, leveraging a representation-based diversity score to capture relationships between data points and retain historical information for comprehensive diversity evaluation. This also allows for flexible combination of diversity and quality scores during data selection and ranking. Extensive experiments demonstrate that PIBE significantly outperforms baselines in InsBank evolution and is able to extract budget-specific subsets, demonstrating its effectiveness and adaptability. [^1] Our code has been released on https://github.com/jiayinlp/InsBank
@@
Jiayi Shi1222Equal contributions.,    Yiwei Li1222Equal contributions.,    Shaoxiong Feng2,    Peiwen Yuan1,    Xinglin Wang1, Yueqi Zhang1,    Chuyi Tan1,    Boyuan Pan2333Corresponding authors.,    Huan Ren2,    Yao Hu 2,    Kan Li1333Corresponding authors., 1 School of Computer Science, Beijing Institute of Technology 2 Xiaohongshu Inc {shijiayi,liyiwei,peiwenyuan,wangxinglin}@bit.edu.cn {shaoxiongfeng2023}@gmail.com   {panboyuan,zengshu,xiahou}@xiaohongshu.com {zhangyq,tanchuyi,likan}@bit.edu.cn
@@ 1 Introduction
Specifically, with the continuous emergence of instruction datasets (The timeline of part instruction datasets is shown in Appendix A), it becomes necessary to regularly update the instruction subset to incorporate the latest advanced instruction data in order to ensure ongoing improvements in the alignment capabilities of LLMs. Simultaneously, the subset size must be controlled to avoid excessive growth that could lead to increased training costs. To address these practical challenges, we propose a novel concept termed InsBank (Instruction Bank). InsBank is designed to support instruction subset evolution with two key properties: (1) To prevent unbounded growth, InsBank maintains a constant size by replacing low-quality old samples with an equal number of high-scoring new ones during evolution. (2) Samples in InsBank are ranked according to their overall scores to enable users to extract subsets that are tailored to specific training budgets, simply by selecting the top-ranked samples. The evolution process of InsBank is illustrated in Figure 1.
As the scale of existing instruction sets continues to grow(Qin et al., 2024a; Longpre et al., 2023; Wang et al., 2023; Xu et al., 2023)—reaching millions or even billions of instances—the cost of exhaustively traversing all candidate data during each InsBank evolution becomes prohibitively high. To address this challenge, we propose Progressive Instruction Bank Evolution (PIBE), a method designed for continuous and efficient selection of the optimal instruction subset. PIBE evolves InsBank in a gradual manner, ensuring long-term efficiency. Unlike the naive approach, it significantly reduces the cost of evolution by excluding previously filtered-out data and focusing solely on newly proposed samples and the current InsBank.
@@
Additionally, the orderliness of InsBank calls for an overall score that integrates both individual quality and diversity signals. While quality scores can be readily obtained through manual or model-based annotation, measuring individual diversity requires global comparisons among candidates. Unfortunately, existing instruction data selection methods struggle to effectively represent and combine quality and diversity for ranking purposes. This challenge is further exacerbated by the absence of historical data, which alters the distribution of candidates and underscores the need to retain historical data distributional information during evolution. Existing diversity-driven data selection methods (Liu et al., 2024; Wu et al., 2023) typically fall into two categories: k-nearest neighbor (k-NN) approaches (Dong et al., 2011) and geometry-based coreset sampling methods (Guo et al., 2022). Both of them rely exclusively on local information from a limited number of neighboring points, which limits their ability to capture global relationships and provide reliable individual diversity scores for ranking. Furthermore, they lack mechanisms to preserve information about previously discarded data, making them ill-suited for progressive selection. Inspired by Affinity Propagation (Frey and Dueck, 2007), we frame InsBank data selection as an exemplar election process, where the representativeness of each data point is quantified through an iterative voting mechanism. The representativeness further serves as the individual diversity score, and the voting results are passed to the next iteration as historical information to preserve the distribution of absent data. Moreover, existing data selection methods either prioritize quality or diversity (Chen et al., 2024), or address them sequentially (Liu et al., 2024), failing to consider both aspects equally. Conversely, our diversity score integrates seamlessly with the quality score, enabling comprehensive and flexible instruction selection and InsBank ranking.
@@
We simulate the instruction set development with five datasets and perform InsBank evolution on them with PIBE and we elaborate on the rationale for selecting these datasets in Appendix J.1. We evaluate the general instruction following capability of fine-tuned models on AlpacaEval (Li et al., 2023b), MT-Bench (Zheng et al., 2023), IFEval (Zhou et al., 2023b), OpenLLM Leaderboard (Beeching et al., 2023) and FollowBench (Jiang et al., 2024). Experimental results show that PIBE outperforms the baselines and successfully evolves the instruction bank in parallel with the development of instruction sets. Besides, analysis on orderliness of InsBank indicates that users can flexibly select a smaller subset based on their budget. Ours contributions can be summarized as follows:
@@ 3 Progressive Instruction Bank Evolution
Figure 2: The framework of PIBE begins by integrating newly proposed candidates with the existing InsBank data and initializing momentum information based on historical records. Then, affinity propagation incorporating the momentum is applied to compute diversity scores. Subsequently, the quality scores obtained via model-based annotation are combined with the diversity scores to produce an individual overall score. Finally, the top-$k$ samples with the highest overall scores are selected to form the evolved InsBank where $k$ is the budget.
@@ 3.1 Gradual Evolution Formulation
In this work, we propose the instruction subset evolution task to build the InsBank. Denoting current available instruction data as $\mathcal{X}_{0}$, the instruction bank $\mathcal{B}_{\pi}^{0,m}$ of size $m$ is initialized through data selection which can be presented as $\mathcal{B}_{\pi}^{0,m}=\pi(\mathcal{X}_{0})$. Then, when new instruction dataset $\mathcal{X}_{1}$ is proposed, $\mathcal{B}_{\pi}^{0,m}$ should evolve itself to adapt to changes in data distribution. The naive manner of InsBank evolution can be represented as $\mathcal{B}_{\pi}^{1,m}=\pi(\mathcal{X}_{0},\mathcal{X}_{1})$ which can be extended to $\mathcal{B}_{\pi}^{t+1,m}=\pi(\mathcal{X}_{0},...,\mathcal{X}_{t},\mathcal{X}_{t+1})$ for future evolution. However, this manner requires substantial storage and computational resources to calculate diversity scores as $t$ continues to increase. To improve the long-term evolution efficiency, we propose a gradual manner where only the newly proposed instruction data $\mathcal{X}_{t+1}$ along with the data participated in last round of evolution $\mathcal{X}_{t}+\mathcal{B}_{\pi}^{t-1,m}$ are involved into the current round of evolution, and the evolution can be represented as $\mathcal{B}_{\pi}^{t+1,m}=\pi(\mathcal{X}_{t+1},\mathcal{X}_{t}+\mathcal{B}_{\pi}^{t-1,m})$.
@@ 3.2 Historical Information Flowing
The structure of $M_{t}$ is depicted in Appendix B. The top-left part of $M_{t}$ contains responsibility values between data in $\mathcal{B}^{t,m}_{\pi}$, taken directly from $H_{t}$. The top-right part represents the suitability of newly proposed candidate data as exemplars for previously selected data, estimated using Eq. 4. Similarly, the bottom-left part represents the suitability of previously selected data as exemplars for newly proposed candidate data, estimated using Eq. 5. The bottom-right section is filled with the median values of the other three sections.
@@ 3.4 Integration of Diversity and Quality
In practice, we observe that further improving quality beyond a certain level can reduce the fine-tuned model’s performance. Additionally, when combining quality and diversity using linear methods, diversity scores often dominate the selection process. This occurs because quality, as a linear score, increases at a constant rate, even when excessively large values provide diminishing benefits. More details can be found in our experimental analysis of score combination (Section 4).
@@
Table 1: Comparison between different methods. For AlpacaEval and MT-Bench, we employ gpt-4o as annotator. The bold text indicates the best results, and the underlined text represents the second-best results. The results of more base models can be found in Appendix J.5.
To address this, we design a nonlinear mapping function for quality scores, shown in Eq. 11. Here, $Q_{p}$ denotes the $p$-th percentile, $r_{l}$ and $r_{h}$ represent the lower and upper percentiles, $S^{\prime}_{q}$ refers to the scaled quality scores, and $\sigma(\cdot)$ is the sigmoid function. The function, illustrated in Figure 7, leverages the sigmoid’s steepness in (2,2)(-2,2) to enhance the distinguishability of scores within $\tau_{l},\tau_{h}]$, while flattening growth for scores above $\tau_{h}$. Data below $\tau_{l}$ are less considered, as such low-quality data are rarely selected into InsBank. Finally, we combine diversity with the nonlinear-mapped quality scores.
@@ 4.1 Experimental Setup
Training and Evaluation In this work, we fine-tune Llama3.2-1B, Llama3.2-3B, Llama3-8B (AI@Meta, 2024), Qwen2.5-7B, Qwen2.5-14B (Qwen Team, 2024) and Mistral-7B (Jiang et al., 2023) on the selected InsBank. Following DEITA (Liu et al., 2024), we set the size of InsBank to 6k for the convenience of subset evolution. We also experiment with InsBank size of 1k and 3k, and the results can be found in Appendix J.4. During training, we further restrict the trainable tokens and the number of conversation turns. We adopt AlpacaEval (Li et al., 2023b), MT-Bench (Zheng et al., 2023) and IFEval (Zhou et al., 2023b) for automatic model alignment performance evaluation. More details about training and evaluation can be found in Appendix C.
@@
  • DEITA Traverse the instruction pool in descending order of quality scores and add a sample to the selected subset if its maximum cosine similarity with existing selected samples is below a threshold (Liu et al., 2024).
@@ 4.4 Analysis
Effectiveness of Diversity and Quality To validate the role of diversity in instruction data selection, we first construct a quality-controlled subset where all data have quality scores within the range of 4.5 to 5.0 (details in Appendix E). Using PIBE, we compute individual diversity scores for the subset, sort the data in descending order, and select the top 6k samples as the most diverse subset and the bottom 6k as the least diverse subset. The distributions of the two subsets are shown in Fig. 6. Before fine-tuning, we restrict the total trainable tokens to 2M. Results in Table 3 indicate that, with comparable quality, models trained on more diverse data achieve better performance.
@@
Table 3: The results of quality-controlled diversity effectiveness experiment. Qua refers to the average quality score, and Div refers to the average diversity score.
When it comes to quality, the improvement from extremely low to high quality is clearly beneficial, as extremely low-quality subsets often contain noisy data, such as irrelevant or incomplete responses. However, is continuously improving quality always effective in the data selection process? To address this, we compare model performance fine-tuned on data selected by the following strategies in the final evolution iteration: (1) Diversity Greedy: selecting data with the highest diversity scores; (2) Quality Greedy: selecting data with the highest quality scores; and (3) PIBE. The results shown in Table 4 reveal a clear trade-off between diversity and quality. A purely greedy approach focusing on either aspect leads to suboptimal outcomes, while a balanced consideration of both proves more effective. This finding aligns with the main experiment results and suggests the existence of a balance point between diversity and quality, which we further investigate through the analysis of score combination.
@@
Analysis of Score Combination We experiment with the different combination methods to explore the contribution of quality and diversity in PIBE.
@@
Table 5: The results of different combination methods. SP- refers to Spearman value, Diff refers to the difference value between SP-Qua and SP-Div.
We first explore the multiplication manner and the addition manner, and the results are reported in Table 5. Overall, regardless of whether addition or multiplication is used as the combination method, the results exhibit a distinct trend of initially increasing and then decreasing as the influence of quality grows (i.e., with the increase of the $\gamma$ value). This finding supports the hypothesis that a balance point exists between diversity and quality.
@@ 5 Related Work
Quality-based Selection Humpback (Li et al., 2023a) selects high-quality samples through an iterative self-curation process where quality predictions are produced by the fine-tuned model of each turn. Recent works typically employ a GPT-model to annotate the data quality. For example, ALPAGASUS (Chen et al., 2024) employs ChatGPT to score the accuracy of instruction data and select data according to a threshold.
@@
Model-specific Importance-based Selection Importance refers to the necessity of adding one sample into training set (Qin et al., 2024a) whose indicator are typically model-specific (Xia et al., 2024; Li et al., 2024a; Hui et al., 2024; Du et al., 2023). However, this work focuses on the general data selection and emphasizes the quality and diversity of selected data.
@@ 6 Conclusion
In this paper, we propose InsBank to address the challenge of evolving instruction subset. PIBE integrates high-quality and representative data into InsBank, striking a balance between data diversity and quality, while maintaining long-term scalability and efficiency. By leveraging a representation-based diversity score with historical information, PIBE flexibly combines diversity and quality for data selection and ranking. Experimental results show PIBE outperforms baselines, providing more optimal and adaptable instruction subsets. The orderliness of InsBank also allows users to extract tailored subsets within budget constraints, supporting cost-effective training and the ongoing refinement of LLMs. This work paves the way for more dynamic and adaptable instruction tuning strategies, enhancing both the efficiency and effectiveness of LLM development over time.
@@
In this work, we focus on evaluating the diversity of individual instruction data and exploring the combination of diversity and quality scores. However, achieving a more precise assessment of data quality remains a valuable direction for future research.